OpenStudy (anonymous):

An airplane flies at an air speed of 300 miles per hour, in the direction toward southwest. There is a head wind of 75 mi/hr in the direction toward due east. (A) Determine the ground speed. (B) Determine the direction of motion of the plane, expressed as an angle counterclockwise from due east.

OpenStudy (shamim):

its a problem of vector

OpenStudy (shamim):

i thnk its a 2 dimensional problem

OpenStudy (shamim):

i think u missed the angle of ur plane velocity 300 mile/hr with the ground

OpenStudy (anonymous):

Choose East as positive x-direction, North is positive y-direction Velocity of the plane ( vector notation ) = -300*sin45 (x) -300*cos45 (y) mi/h (x) : noted as x-direction, (y) : noted as y-direction, - negative sign indicates negative directions. Velovity of the wind = 75 (x) mi/h (A) the ground speed = Square root of (V(airplane)^2 + V(wind)^2) = Sqr of ((-212.13 + 75)^2 (x) + (-212.13)^2 (y)) = 252.6 mi/h (B) To determine the direction counterclockwise due to East, use tangent equation to fine the angle due to East. Ground velocity (in vector notation)= -137.3 (x) - 212.13 (y) Angle & ---> Tan(&) = y/x =1.545 -----> & = Tan^-1(1.545)=57 degree This is the angle between South and West, but we want to find the angle due to East. So the angle we want = 180-57 =123 degree due east